A particle moving with acceleration 6 m/s2 along x-axis covers 30 m in 5th second. The distance covered by it in 4th second is
 

Grade:11

1 Answers

Arun
6 years ago
As we know that distance covered in nth second
Sn = u +(1/2) * f* (2n-1)
Hence
30 = u + (6/2) * (10-1)
u = 30 -27 = 3 m/sec
 
Hence distance covered in 4th second
S4 = 3 + (1/2)* 6* (9-1)
 = 3 + 24
= 27 metres